OpenStudy (anonymous):

Please help on 3 questions! Fan and Medal! <3 :)

OpenStudy (anonymous):

1.

OpenStudy (anonymous):

2. Use a related equation that has the variable on one side. Then simplify the other side. p + 6 = 18 A. p = 3 B. p = 12 C. p = 24 D. p = 108

OpenStudy (anonymous):

3. What is the value of the variable? 13c = 39 c =

OpenStudy (anonymous):

3 c= 3? c=39/13 =3

OpenStudy (anonymous):

so what nukber is it XD

OpenStudy (anonymous):

C=3

OpenStudy (anonymous):



OpenStudy (anonymous):

:D